# Condensates in aggregation processes with input: A new power law and intermittency

### v2

@@ Abstract
Large-mass condensates, which coexist with a power-law-decaying distribution in the one-dimensional Takayasu model of mass aggregation with input, were recently found in numerical simulations. Here we establish the occurrence of condensates by analyzing exact recursions for finite systems, and further show that they have a strong effect on the properties of the system. In the steady state of a large but finite system, there is a single condensate, whose random movement through the system leads to a reorganization of the mass profile on a macroscopic scale. A scaling analysis of the mean mass and standard deviation at different distances from the condensate leads to the surprising conclusion that the mass distribution on a macroscopic number of sites around the condensate follows a power-law decay with an exponent 5/3, while further-away sites show the customary Takayasu exponent 4/3, with a crossover in between. Finally, the exit of condensates from a system with open boundaries has a strong effect on the temporal fluctuations of the total mass in the steady state. Their departure is followed by a build-up of mass and subsequent departures, leading to strong intermittency, established through a divergence of the flatness as the scaled time approaches zero.
@@ I Introduction
Aggregation phenomena are important in many areas of the natural sciences, cutting across different fields and length scales, from clouds and river networks on the one hand, to colloid formation and biomolecular organization on the other Friedlander and Smoke (2000); Scheidegger (1967); Smoluchowski (1916); Brangwynne et al. (2011); Berry et al. (2018); Lee et al. (2023); Arsenadze et al. (2024). The subject, which began with the pioneering studies of colloidal clusters by Smoluchowski Smoluchowski (1916), continues to find new domains of applicability, most recently in aspects of biomolecular organization Brangwynne et al. (2011); Arsenadze et al. (2024); Berry et al. (2018); Lee et al. (2023).
@@
The kinetics of aggregation involves the coalescence and breakup of clusters with different numbers of particles, thus different masses. The resulting time-dependent mass distribution often shows interesting features. First, the distribution may be broad, characterized by a power law, signalling scale invariance arising from critical correlations in the system. Secondly, spatially well-separated condensates may form, each one being a single cluster which holds a very large number of constituent particles. It is natural to ask: Under what conditions are these features manifest???? Can they arise together, and if so, under what circumstances????
@@
Much can be learned from the analysis of simple models whose definitions include only the essential kinetic processes, but whose emergent behaviour is nontrivial. The Takayasu model of aggregation is a prime example Takayasu et al. (1988); Privman (1997); Takayasu et al. (1991). Its moves include diffusion of particle clusters, followed by coalescence when two clusters meet. Further, there is an input of particles at a steady rate. The appeal of the model stems from its simplicity and the fact that as time passes, the distribution of particle numbers approaches a power-law scaling function of mass and time, indicating a self-organized critical (SOC) state. This can be established analytically within mean field theory and through an exact recursive calculation in 1D Takayasu et al. (1988); Takayasu (1989); Privman (1997); Huber (1991). In addition, spatiotemporal correlation functions can be calculated exactly Rajesh and Majumdar (2000). Interestingly, the Takayasu model has exact correspondences to several other models, for instance the Scheidegger model of river basins Scheidegger (1967); Rajesh and Majumdar (2000), the voter model Liggett (2013) and models of stress propagation in bead packs Coppersmith et al. (1996). Further, experiments indicate that nucleoli in certain living cells share the Takayasu moves of diffusion, coalescence and input, and show a broad power-law distribution of sizes Brangwynne et al. (2011); Berry et al. (2018); Lee et al. (2023).
@@
Whereas the power law scaling is well established and known since the inception of the Takayasu model, it is only recently that it was realized that the model also exhibits condensates or extremely large-mass clusters Das and Barma (2023) at both intermediate and long times. Condensates are known to arise in aggregation models with mass conservation, where their formation is akin to Bose-Einstein condensation Evans and Hanney (2005); Majumdar et al. (1998, 2000a); Das et al. (2015); Iyer et al. (2023). But in the Takayasu model, where the mass is strongly non-conserved owing to the steady input, the mechanism is quite different. It is an effect arising from the influence of the maximal term in slowly decaying power-law distributions Darling (1952); Feller (1991). Numerical simulations show that due to this maximal term effect, beyond the power law, the probability distribution of masses in the Takayasu model exhibits a strong and distinctive condensate hump, which is part of the scaling function Das and Barma (2023), confirmed by the numerically exact recursive results reported below. A time-dependent length scale $\mathcal{L}(t)$ determines condensate masses and the spacing between them. Since the underlying process is diffusive, $\mathcal{L}(t)$ is proportional to $t^{1/2}$. This implies that as $t$ increases, the number of condensates falls while the size of each increases, as in phase ordering dynamics Iyer et al. (2023); Bray (1994). This continues till $\mathcal{L}(t)$ reaches the system size $L$, beyond which a long time state with statistically invariant properties sets in Das and Barma (2023).
@@
Condensates have a strong effect on the properties of the Takayasu model. In this paper, we first establish their occurrence in finite systems by an exact calculation of the probability distribution of mass as a function of time (see Fig. 1a) through recurrence relations of the mass held in a stretch of lattice sites Takayasu et al. (1988); Privman (1997); Huber (1991). Further, we investigate the effect of condensates on the mass profile in the system, recognizing that a condensate functions like an efficient forager which ingests mass from neighboring regions. The result is surprising: with periodic boundary conditions, the mean mass profile seen by the condensate is strongly influenced on a macroscopic scale. The mean and standard deviation are both scaling functions of $r/L$ where $r$ is the distance from the condensate. As shown below, the results imply that the distribution of masses on sites close to the condensate follows a new power law $\sim m^{-5/3}$ as opposed to the well-known Takayasu power $\sim m^{-4/3}$, which holds on distant sites. Finally, we investigate the time variation of the mass of a system with free boundary conditions and find that the build-up and departure of condensates leads to enormous fluctuations of the total mass, characterized by strong intermittency, reminiscent of dragon king events Sornette and Ouillon (2012); Mikaberidze et al. (2023).
@@ II The Model and earlier work
The dynamics of the system is thus described by the stochastic equation
@@
Earlier, it was found that in the limit $L\rightarrow\infty$, followed by $t\rightarrow\infty$, the mass distribution follows a power-law decay $p(m)\sim m^{-4/3}$ Takayasu (1989). If $t$ is finite, there is a time-dependent length proportional to $t^{1/2}$ within which masses are correlated and beyond which they are not. An exact calculation shows that the equal-time two point correlation function follows a scaling form $\langle m(i,t)m(i+r,t)\rangle_{c}=t^{2}f(r/\mathcal{L}(t))$, with $\mathcal{L}(t)=t^{1/2}$, in the limit $r\rightarrow\infty,t\rightarrow\infty$ with $r/\mathcal{L}(t)$ finite Rajesh and Majumdar (2000).
Recent numerical simulations have revealed the presence of condensates in the system Das and Barma (2023), which lead to a characteristic hump in the mass distribution at finite times. The condensate accumulates a significant portion ($\sim 80\%$) of the total mass in the system, and in the long-time limit, the system contains a single condensate with mass of order $O(L^{3})$.
@@ III The maximal term and condensates
Consider $n$ random variables $m_{1},m_{2},....m_{n}$. Suppose they are independent and identically distributed with the probability distribution function $p(m)\sim m^{-(1+\alpha)}$ with $\alpha<1$. Let $S_{n}$ and $M_{n}$ denote their sum and maximum respectively: $S_{n}=\sum_{i=1}^{n}m_{i}$ and $M_{n}=\max(m_{1},m_{2},...,m_{n})$. The variables $S_{n}$ and $M_{n}$ are strongly correlated and satisfy Darling (1952); Feller (1991)
@@
The result Eq. (2) holds for independent, identically distributed variables, and the question arises whether a similar mechanism for condensate formation operates also when variables are correlated. For the Takayasu model, the answer is in the affirmative despite correlations between masses on different sites. It seems that the existence of the condensate rests on the occurrence of the slow power law $p(m)\sim m^{-4/3}$, while the correlations between masses have the relatively minor effect of changing the asymptotic value of $\langle S_{n}/M_{n}\rangle$ from $3/2$ (in the absence of correlations) to $\simeq 1.41$ (obtained by numerically averaging $S_{n}/M_{n}$ in subsystems of size $\mathcal{L}(t)=t^{1/2}$ over many histories.)
@@ IV Exact solution
Following Takayasu (1989), we consider the mass contained in a stretch of $r$ consecutive sites at any instant:
@@
where the average is over all histories. Equations (1) and (3), lead to the recursion relation Takayasu (1989)
@@
with the initial condition $Z_{r}(q,0)=\exp{(iqr)}$, and the boundary conditions $Z_{0}(q,t)=1$ and $Z_{L}(q,t)=\exp(iqL(t+1))$. Here, the prefactor $\phi(q)\equiv\langle\exp\left\{iqI_{i}(t)\right\}\rangle$. Earlier, Takayasu et al. Takayasu (1989); Takayasu et al. (1991) and Huber Huber (1991) obtained the $L=\infty$, $t=\infty$ mass distribution to leading order in $m$. They proceeded by setting $Z_{r}(q,t+1)=Z_{r}(q,t)$ in Eq. (5) and expanding $\phi(q)$ as
@@
We remark that the condensates did not appear in the treatments of Takayasu (1989); Takayasu et al. (1991); Huber (1991), as they are missed in the limit $t\rightarrow\infty$ in an infinite system. We obtain exact results for finite $L$ and $t$ below and show the existence of condensates, corroborating the conclusions of Das and Barma (2023), which were based on numerical simulations.
To this end, we first iterate Equation (5) $t$ times for a given $L$, with the help of a computer, to obtain $Z_{1}(q,t)$, and then perform an inverse Fourier transform on $Z_{1}(q,t)$ yielding
@@
In a system of size $L$, there is a characteristic time $t^{*}=L^{2}$ over which mass fluctuations spread diffusively across the system. For $t\ll t^{*}$, fluctuations spread across a length $\mathcal{L}(t)=\sqrt{t}\ll L$, whereas for $t$ exceeding $t^{*}$, the system reaches a ‘steady state’ in a sense to be made precise below. Below, we discuss our results for $p(m,t)$ separately for the two regimes: (a) $t\ll t^{*}$ and (b) $t\gg t^{*}$. We will find that condensates show up in both the regimes — as a hump in the former regime and as a sharper peak in the latter. We restrict ourselves to relatively small values of $L$, namely, $L=8$ and $16$, as they suffice to show all the relevant features of $p(m,t)$ and as going to higher values of $L$ is computationally costly.
(a) To study $p(m,t)$ in the regime $t\ll t^{*}$, we examine the cases $t=10$ and $15$ for $L=8$ and $16$ shown in Fig. 1a: the distribution is characterized by the power-law $p(m)\sim m^{-4/3}$ followed by a hump. The hump corresponds to the presence of large condensates in the system, with typically one condensate in each region of length $\mathcal{L}(t)$ Das and Barma (2023). The distribution is cut-off beyond the hump at $m=(t+1)L$, which is the total mass in the system at time $t$. For both values of $L$, the distributions $p(m,t)$ are identical except for the cut-off, a property which we expect will continue to hold for any $L$. For different values of $t$, on scaling $m$ by $t^{-3/2}$ and $p(m,t)$ by $t^{2}$, for $t\ll t^{*}$, the curves collapse (see Fig. 1b), implying the scaling form $p(m,t)\approx t^{-2}f(u)$, where $u=mt^{-3/2}$. Note that the condensate hump is part of the scaling function. The region $u<u_{0}\simeq 0.6$ shows the aforementioned power-law decay, while the condensate hump appears for $u>u_{0}$, with $f(u)$ decaying exponentially as $u\rightarrow\infty$ Das and Barma (2023).
@@
(b) For a fixed $t$ in the regime $t\gg t^{*}$, the distribution follows $p(m,t)\sim m^{-4/3}$ for $m\ll m^{*}$ and $p(m)\sim\exp(-m/m^{*})$ for $m\gg m^{*}$, where $m^{*}=L^{3}+L$ is a characteristic mass equal to the total mass in the system up to time $t^{*}$. Towards the end of this exponential tail the distribution exhibits a prominent peak, representing a single large condensate. The area under the peak is $1/L$, the probability to find the condensate at a randomly chosen site. In the steady state, the condensate peak in the distribution is seen to follow $p(m,t)\approx H(m-Lt)$ (see Fig. 1c), showing that once $t$ exceeds $t^{*}$, practically all the mass injected into the system is eventually absorbed by the condensate. Concomitantly, the mass distribution of the rest of the system, which corresponds to the portion before the condensate peak, assumes a time-independent form. Thus, we shall refer to the $t\gg t^{*}$ regime as ‘steady state’, despite the fact that the condensate mass keeps growing.
@@ V.1 Analytic form for $\mu(r,L)$
Consider a system of size $L$ with periodic boundary conditions following the Takayasu dynamics (see Eq.(1) in the main text). Let $\mu(r,L)$ and $J(r,L)$ denote the mean mass and the mean mass current, respectively, at position $r$ with respect to the location of the condensate at steady state. As the masses follow symmetric diffusive dynamics, we write
$J(r,L)=D[\mu(r,L)-\mu(r-1,L)-(\mu(r+1,L)-\mu(r,L))],$ (8)
@@
$J=-D\partial^{2}\mu/\partial r^{2}.$ (9)
Evidently, at steady state, $J(r,L)$ must satisfy the continuity equation
@@
Figure 4: (a) Scaling of the CCDF for the masses on sites next to the condensate with $L$ for $L=10$, $20$, $50$, and, $100$ at steady state. Inset: Unscaled plots. (b) Scaling of the CCDF for masses on the sites next to the condensate with $t$ in the regime $t\ll t^{*}$ for $L=1200$. The inset shows CCDF vs $m$ plotted at different times for masses on the sites next to the condensate and masses in the bulk along with the lines $m^{-2/3}$ and $m^{-1/3}$, respectively.
@@
The CCDF of masses relative to the condensate depicts a clear crossover as $r/L$ varies, with the exponent $\theta$ changing from $4/3$ to $5/3$. When $r/L$ is small, the $\theta=4/3$ region is small, and the $\theta=5/3$ region dominates. As $r/L$ increases, the $\theta=4/3$ region continues to expand until it prevails everywhere. Thus, a macroscopic number of sites exhibit a power-law $P(m)\sim m^{-5/3}$ in steady state. We remark that the same power law has been found in the case of the charge-injection model Takayasu (1989). When positive and negative charges are input with equal probability, we have $\langle I\rangle=0$. Then the leading behavior of $p(m)$ with $m$, where $m$ now represents the net charge at any site, is determined by the $\langle I^{2}\rangle$ term in Eq. (6), yielding the power-law $p(m)\sim m^{-5/3}$ to leading order in $m$. For mass-injection, we have $\langle I\rangle\neq 0$; however, the mass distribution in a finite fraction of the sites still shows the power-law $P(m)\sim m^{-5/3}$. It would be interesting if a connection could be drawn between the two cases.
[…]
@@
These steady state results also have interesting implications for the regime $t\ll t^{*}$. As diffusive correlations in the system lead to the formation of subsystems of size $\mathcal{L}(t)$, this length scale effectively replaces the system size $L$ in this regime. The main plot of Fig. 4b shows the CCDF of masses next to the condensate scaled using $\mathcal{L}(t)$ in place of $L$. The inset in Fig. 4b shows the unscaled CCDFs in the bulk and adjacent to the condensate at different times, highlighting the presence of two distinct power law behaviors.

### v3

# Condensate-induced organization of the mass profile and emergent power laws in the Takayasu aggregation model

@@ Abstract
Large-mass condensates, which coexist with a power-law-decaying distribution in the one-dimensional Takayasu model of mass aggregation with input, were recently found in numerical simulations. Here, we establish the occurrence of condensates by analyzing exact recursions for finite systems and further show that they have a strong effect on the properties of the system. In the steady state of a large but finite system, there is a single condensate, whose random movement through the system leads to a reorganization of the mass profile on a macroscopic scale. A scaling analysis of the mean mass and standard deviation at different distances from the condensate leads to the surprising conclusion that the mass distribution on a macroscopic number of sites around the condensate follows a power-law decay with an exponent 5/3, while farther-away sites show the customary Takayasu exponent 4/3, with a crossover in between. Finally, the exit of condensates from a system with open boundaries has a strong effect on the temporal fluctuations of the total mass in the steady state. Their departure is followed by a buildup of mass and subsequent departures, leading to strong intermittency, established through a divergence of the flatness as the scaled time approaches zero.
@@ I Introduction
Aggregation phenomena are important in many areas of the natural sciences, cutting across different fields and length scales, from clouds and river networks on the one hand, to colloid formation and biomolecular organization on the other [1, 2, 3, 4, 5, 6, 7]. The subject, which began with the pioneering studies of colloidal clusters by Smoluchowski [3], continues to find new domains of applicability, most recently in aspects of biomolecular organization [4, 7, 5, 6].
@@
The kinetics of aggregation involves the coalescence and breakup of clusters with different numbers of particles and, thus different masses. The resulting time-dependent mass distribution often shows interesting features. First, the distribution may be broad, characterized by a power law, signaling scale invariance arising from critical correlations in the system. Secondly, spatially well-separated condensates may form, each one being a single cluster that holds a very large number of constituent particles. It is natural to ask the following questions: Under what conditions do these features manifest???? Can they arise together, and if so, under what circumstances????
@@
Much can be learned from the analysis of simple models whose definitions include only the essential kinetic processes, but whose emergent behavior is nontrivial. The Takayasu model of aggregation is a prime example [8, 9, 10]. Its moves include diffusion of particle clusters, followed by coalescence when two clusters meet. Further, there is an input of particles at a steady rate. The appeal of the model stems from its simplicity and the fact that as time passes, the distribution of particle numbers approaches a power-law scaling function of mass and time, indicating a self-organized critical (SOC) state. This can be established analytically within mean field theory and through an exact recursive calculation in 1D [8, 11, 9, 12]. In addition, spatiotemporal correlation functions can be calculated exactly [13]. Interestingly, the Takayasu model has exact correspondences to several other models, for instance the Scheidegger model of river basins [2, 13], the voter model [14] and models of stress propagation in bead packs [15]. Further, experiments indicate that nucleoli in certain living cells share the Takayasu moves of diffusion, coalescence and input, and show a broad power-law distribution of sizes [4, 5, 6].
@@
Whereas the power law scaling is well established and known since the inception of the Takayasu model, it is only recently that it was realized that the model also exhibits condensates or extremely large-mass clusters [16] at both intermediate and long times. Condensates are known to arise in aggregation models with mass conservation, where their formation is akin to Bose-Einstein condensation [17, 18, 19, 20, 21]. But in the Takayasu model, where the mass is strongly non-conserved owing to the steady input, the mechanism is quite different. It is an effect arising from the influence of the maximal term in slowly decaying power-law distributions [22, 23]. Numerical simulations show that due to this maximal term effect, beyond the power law, the probability distribution of masses in the Takayasu model exhibits a strong and distinctive condensate hump, which is part of the scaling function [16], confirmed by the numerically exact recursive results reported below. A time-dependent length scale $\mathcal{L}(t)$ determines condensate masses and the spacing between them. Since the underlying process is diffusive, $\mathcal{L}(t)$ is proportional to $t^{1/2}$. This implies that as $t$ increases, the number of condensates falls while the size of each increases, as in phase ordering dynamics [21, 24]. This continues until $\mathcal{L}(t)$ reaches the system size $L$, beyond which a long time state with statistically invariant properties sets in [16].
@@
Condensates have a strong effect on the properties of the Takayasu model. In this paper, we first establish their occurrence in finite systems by an exact calculation of the probability distribution of mass as a function of time (see Fig. 1a) through recurrence relations of the mass held in a stretch of lattice sites [8, 9, 12]. Further, we investigate the effect of condensates on the mass profile in the system, recognizing that a condensate functions like an efficient forager which ingests mass from neighboring regions. The result is surprising: with periodic boundary conditions, the mean mass profile seen by the condensate is strongly influenced on a macroscopic scale. The mean and standard deviation are both scaling functions of $r/L$ where $r$ is the distance from the condensate. As shown below, the results imply that the distribution of masses on sites close to the condensate follows a new power law $\sim m^{-5/3}$ as opposed to the well-known Takayasu power $\sim m^{-4/3}$, which holds on distant sites. Finally, we investigate the time variation of the mass of a system with free boundary conditions and find that the build-up and departure of condensates leads to enormous fluctuations of the total mass, characterized by strong intermittency, reminiscent of dragon king events [25, 26].
@@ II The Model and earlier work
The dynamics of the system are thus described by the stochastic equation
@@
Earlier, it was found that in the limit $L\rightarrow\infty$, followed by $t\rightarrow\infty$, the mass distribution follows a power-law decay $p(m)\sim m^{-4/3}$ [11]. If $t$ is finite, there is a time-dependent length proportional to $t^{1/2}$ within which masses are correlated and beyond which they are not. An exact calculation shows that the equal-time two-point correlation function follows a scaling form $\langle m(i,t)m(i+r,t)\rangle_{c}=t^{2}f(r/\mathcal{L}(t))$, with $\mathcal{L}(t)=t^{1/2}$, in the limit $r\rightarrow\infty,t\rightarrow\infty$ with $r/\mathcal{L}(t)$ finite [13].
Recent numerical simulations have revealed the presence of condensates in the system [16], which lead to a characteristic hump in the mass distribution at finite times. The condensate accumulates a significant portion ($\sim 80\%$) of the total mass in the system, and in the long-time limit, the system contains a single condensate with mass of order $O(L^{3})$.
@@ III The maximal term and condensates
Consider $n$ random variables $m_{1},m_{2},....m_{n}$. Suppose they are independent and identically distributed with the probability distribution function $p(m)\sim m^{-(1+\alpha)}$ with $\alpha<1$. Let $S_{n}$ and $M_{n}$ denote their sum and maximum respectively: $S_{n}=\sum_{i=1}^{n}m_{i}$ and $M_{n}=\max(m_{1},m_{2},...,m_{n})$. The variables $S_{n}$ and $M_{n}$ are strongly correlated and satisfy [22, 23]
@@
Equation (2) holds for independent, identically distributed variables, and the question arises whether a similar mechanism for condensate formation operates also when variables are correlated. For the Takayasu model, the answer is in the affirmative despite correlations between masses on different sites. It seems that the existence of the condensate rests on the occurrence of the slow power law $p(m)\sim m^{-4/3}$, while the correlations between masses have the relatively minor effect of changing the asymptotic value of $\langle S_{n}/M_{n}\rangle$ from $3/2$ (in the absence of correlations) to $\simeq 1.41$ (obtained by numerically averaging $S_{n}/M_{n}$ in subsystems of size $\mathcal{L}(t)=t^{1/2}$ over many histories.)
An interpretation of the condensate mass in terms of Lévy distributions is possible even if $\alpha>1$ [27], in which case $S_{n}$ would be proportional to $n$. The interesting point here is that Eq. (2) shows that the interpretation holds even if $\alpha<1$ and $S_{n}$ grows as $n^{1/\alpha}$.
@@ IV Exact solution
Following [11], we consider the mass contained in a stretch of $r$ consecutive sites at any instant:
@@
where the average is over all histories. Equations (1) and (3), lead to the recursion relation [11]
@@
with the initial condition $Z_{r}(q,0)=\exp{(iqr)}$, and the boundary conditions $Z_{0}(q,t)=1$ and $Z_{L}(q,t)=\exp(iqL(t+1))$. Here, the prefactor $\phi(q)\equiv\langle\exp\left\{iqI_{i}(t)\right\}\rangle$. Earlier, Takayasu et al. [11, 10] and Huber [12] obtained the $L=\infty$, $t=\infty$ mass distribution to leading order in $m$. They proceeded by setting $Z_{r}(q,t+1)=Z_{r}(q,t)$ in Eq. (5) and expanding $\phi(q)$ as
@@
We remark that the condensates did not appear in the treatments of [11, 10, 12], as they are missed in the limit $t\rightarrow\infty$ in an infinite system. We obtain exact results for finite $L$ and $t$ below and show the existence of condensates, corroborating the conclusions of [16], which were based on numerical simulations.
To this end, we first iterate Eq. (5) $t$ times for a given $L$, with the help of a computer, to obtain $Z_{1}(q,t)$, and then perform an inverse Fourier transform on $Z_{1}(q,t)$ yielding
@@
In a system of size $L$, there is a characteristic time $t^{*}=L^{2}$ over which mass fluctuations spread diffusively across the system. For $t\ll t^{*}$, fluctuations spread across a length $\mathcal{L}(t)=\sqrt{t}\ll L$, whereas for $t$ exceeding $t^{*}$, the system reaches a steady state in a sense to be made precise below. Below, we discuss our results for $p(m,t)$ separately for the two regimes: (a) $t\ll t^{*}$ and (b) $t\gg t^{*}$. We will find that condensates show up in both the regimes — as a hump in the former regime and as a sharper peak in the latter. We restrict ourselves to relatively small values of $L$, namely, $L=8$ and $16$, as they suffice to show all the relevant features of $p(m,t)$ and as going to higher values of $L$ is computationally costly.
(a) To study $p(m,t)$ in the regime $t\ll t^{*}$, we examine the cases $t=10$ and $15$ for $L=8$ and $16$ shown in Fig. 1a: the distribution is characterized by the power-law $p(m)\sim m^{-4/3}$ followed by a hump. The hump corresponds to the presence of large condensates in the system, with typically one condensate in each region of length $\mathcal{L}(t)$ [16]. The distribution is cutoff beyond the hump at $m=(t+1)L$, which is the total mass in the system at time $t$. For both values of $L$, the distributions $p(m,t)$ are identical except for the cutoff, a property that we expect will continue to hold for any $L$. For different values of $t$, on scaling $m$ by $t^{-3/2}$ and $p(m,t)$ by $t^{2}$, for $t\ll t^{*}$, the curves collapse (see Fig. 1b), implying the scaling form $p(m,t)\approx t^{-2}f(u)$, where $u=mt^{-3/2}$. Note that the condensate hump is part of the scaling function. The region $u<u_{0}\simeq 0.6$ shows the aforementioned power-law decay, while the condensate hump appears for $u>u_{0}$, with $f(u)$ decaying exponentially as $u\rightarrow\infty$ [16].
@@
(b) For a fixed $t$ in the regime $t\gg t^{*}$, the distribution follows $p(m,t)\sim m^{-4/3}$ for $m\ll m^{*}$ and $p(m)\sim\exp(-m/m^{*})$ for $m\gg m^{*}$, where $m^{*}=L^{3}+L$ is a characteristic mass equal to the total mass in the system up to time $t^{*}$. Toward the end of this exponential tail the distribution exhibits a prominent peak, representing a single large condensate [see Fig. 1a, L = 8, 16; t = 500]. The area under the peak is $1/L$, the probability to find the condensate at a randomly chosen site. In the steady state, the condensate peak in the distribution is seen to follow $p(m,t)\approx H(m-Lt)$ (see Fig. 1c), showing that once $t$ exceeds $t^{*}$, practically all the mass injected into the system is eventually absorbed by the condensate. Concomitantly, the mass distribution of the rest of the system, which corresponds to the portion before the condensate peak, assumes a time-independent form. Thus, we shall refer to the $t\gg t^{*}$ regime as ‘steady state’, despite the fact that the condensate mass keeps growing.
@@ V.1 Analytic form for $\mu(r,L)$
Consider a system of size $L$ with periodic boundary conditions following the Takayasu dynamics (see Eq.(1)). Let $\mu(r,L)$ and $J(r,L)$ denote the mean mass and the mean mass current, respectively, at position $r$ with respect to the location of the condensate at steady state. As the masses follow symmetric diffusive dynamics, we write
$J(r,L)=-D[\mu(r,L)-\mu(r-1,L)],$ (8)
@@
$J=-D\partial\mu/\partial r.$ (9)
Evidently, in steady state, $J(r,L)$ must satisfy the continuity equation with injection
@@
Figure 4: (a) Scaling of the complementary cumulative distribution function (CCDF) for the masses on the sites next to the condensate with $L$ for $L=10$, $20$, $50$, and, $100$ at steady state. Inset: Unscaled plots. (b) Scaling of the CCDF for masses on the sites next to the condensate with $t$ in the regime $t\ll t^{*}$ for $L=1200$. The inset shows CCDF vs $m$ plotted at different times for masses on the sites next to the condensate and masses in the bulk along with the lines $m^{-2/3}$ and $m^{-1/3}$, respectively.
@@
The CCDF of masses relative to the condensate depicts a clear crossover as $r/L$ varies, with the exponent $\theta$ changing from $4/3$ to $5/3$ (see Fig. 5). When $r/L$ is small, the $\theta=4/3$ region is small, and the $\theta=5/3$ region dominates. As $r/L$ increases, the $\theta=4/3$ region continues to expand until it prevails everywhere. Thus, a macroscopic number of sites exhibit a power-law $P(m)\sim m^{-5/3}$ in steady state. We remark that the same power law has been found in the case of the charge-injection model [11]. When positive and negative charges are input with equal probability, we have $\langle I\rangle=0$. Then the leading behavior of $p(m)$ with $m$, where $m$ now represents the net charge at any site, is determined by the $\langle I^{2}\rangle$ term in Eq. (6), yielding the power-law $p(m)\sim m^{-5/3}$ to leading order in $m$. For mass injection, we have $\langle I\rangle\neq 0$; however, the mass distribution in a finite fraction of the sites still shows the power-law $P(m)\sim m^{-5/3}$. It would be interesting if a connection could be drawn between the two cases.
@@
These steady state results also have interesting implications for the regime $t\ll t^{*}$.Due to diffusive correlations, the system can be thought of as comprised of subsystems of size $\mathcal{L}(t)$. This length scale effectively replaces the system size $L$ in this regime. The main plot of Fig. 4b shows the CCDF of masses next to the condensate scaled using $\mathcal{L}(t)$ in place of $L$. The inset in Fig. 4b shows the unscaled CCDFs in the bulk and adjacent to the condensate at different times, highlighting the presence of two distinct power law behaviors.